# Improving causal inference in interrupted time series analysis: the triple difference design

### 18 Mar 2026 v1

@@ 5 Conclusion
The DDD-ITSA design offers a valuable addition to the health policy researcher’s toolkit, extending the logic of multiple-group interrupted time series to address remaining sources of confounding through the inclusion of a second control group. By differencing across time, groups, and subgroups, DDD-ITSA enhances the credibility of causal claims in observational interrupted time-series settings when suitable control series are available.
Supplementary Information
The Appendix contains Stata code for replicating the Example.
Authors’ contributions
AL conceived the study and its design, conducted all analyses, wrote the manuscript and takes public responsibility for its content.
Funding
There was no funding associated with this work.
Declarations
Ethics approval and consent to participate: Not applicable.
Consent for publication: Not applicable.
Competing interests
The author declares no competing interests.

### 25 Apr 2026 v2

@@ 5 Conclusion
The DDD-ITSA design offers a valuable addition to the health policy researcher’s toolkit, extending the logic of multiple-group interrupted time series to address remaining sources of confounding through the inclusion of a second control group. By formalizing the model, clarifying parameter interpretation, and providing practical implementation guidance, this paper aims to facilitate appropriate use of the design. The updated itsa package for Stata makes DDD-ITSA accessible for routine application, enabling researchers to improve causal inference when the identifying assumptions of the design are plausible and supported by substantive knowledge and empirical assessment. As with all quasi-experimental methods, the validity of DDD-ITSA estimates ultimately depends on the plausibility of its identifying assumptions in the specific context of application, underscoring the importance of substantive knowledge, transparent reporting, and critical evaluation of evidence.